By using single layer perceptron algorithm classify the following patterns. Where X1 and X2

are inputs Y is the output. Learning rate is 0.5 , output is 1 if weighted sum >=0.5 otherwise

0 and weights w1 =0.2 and w2 = 0.3 [5]

X1 X2 Y

0 0 0

0 1 1

1 0 1

1 1 1


1
Expert's answer
2021-07-12T17:18:53-0400

For the linearly separable,

"w_1x_1+w_2x_2+\eta=0"

"\eta = 0.5"

"w_1=0.2"

"w_2=0.3"

"x_2= \frac{w_1}{w_2}x_1-\frac{\eta}{w_2}"

Now, substituting the values,

"x_2= \frac{-0.2}{0.3}x_1-\frac{0.5}{0.3}"